The potential difference between two parallel plates of a condenser which are 4 mm apart is 800 V. (a) What is the potential gradient? (b) What is the force on a proton which is midway between the plates?



With complete Solution Thank you

1
Expert's answer
2022-02-09T06:29:35-0500

Given:

"d=\rm 4 \: mm=0.004\: m"

"V=800\:\rm V"


(a) the potential gradient

"E=\frac{V}{d}=\rm \frac{800\: V}{0.004\: m}=2*10^5\: V/m"

(b) the force on a proton which is midway between the plates


"F=qE\\
=\rm 1.6*10^{-19}\: C*2*10^5\: V/m=3.2*10^{-14}\: N"